OpenStudy (anonymous):

Prove that each statement is true for all positive integers. What are the fist few steps? 1 * 3 + 3 * 5 + 5 * 7 + ... + (2n-1)(2n+1) = n(4n^2 + 6n - 1) / 3

OpenStudy (anonymous):

induction proof?

OpenStudy (anonymous):

what kind of proof are you needing? i think you have to do it by induction

OpenStudy (anonymous):

\[\sum _{k=1}^n \left(4 k^2-1\right)= \frac{1}{3} n \left(4 n^2+6 n-1\right) \] Notice that it is true for n=1. Suppose that it is true for for the sum up to n and prove it for the sum up to n+1 which means suppose that \[\sum _{k=1}^n \left(4 k^2-1\right)= \frac{1}{3} n \left(4 n^2+6 n-1\right) \] and show that \[\sum _{k=1}^{n+1} \left(4 k^2-1\right)= \frac{1}{3} (n+1) \left(4 (n+1)^2+6 ( n+1)-1\right) \]\\ \]